Help!

nment FULL SCREEN PRINTER VERSION BACK Question 5 An article in the Australian Journal of Agricultural Research, Non-Starch Polysaccharides and Broiler Performance on Diets Containing Soyabean Meal as the Sole Protein Concentrate (1993, Vol. 44, No. 8, pp. 1483-1499) determined the essential amino acid (Lysine) composition level of soybean meals are as shown below (g/kg): 22.2 24.7 20.9 26,.0 27.0 24.8 26.5 23.8 25.6 23.9 Round your answers to 2 decimal places. (a) Construct a 99% two-sided confidence interval for σ (b) Calculate a 99% lower confidence bound for (c) Calculate a 90% lower confidence bound for σ
Answer #1

From the given data we have

s = 1.9126
s^2 = 3.6582

a)
chi square value for lower bound = 23.5894
chi square value for upper bound = 1.7349

CI = ((10 - 1)*3.6582/23.5894, (10 - 1)*3.6582/1.7349)
= (1.3956, 18.9762)

b)
(10 - 1)*3.6582/23.5894 = 1.3956

1.3956 <= sigma^2

c)
chi square value for lower bound = 16.9190

(10 - 1)*3.6582/16.9190 = 1.9459

1.9459 <= sigma^2